Hi can please help me to solve this problem? Thank you for your time and efforts. Here is the question? The seawater Orthlieb Pool in Casablanca Morocco, is the largest swimming pool in the world. With a perimeter of 1,110 meters, this rectangular-shaped pool is 30 meters longer than 6 times its width. Find its dimensions?
Answer by mananth(16946)

width be x
length = 6x+30
Perimeter = 1110
Perimeter = 2(l+w)
1110=2(x+6x+30)
1110= 2(7x+30)
1110/2 = 7x+30
555=7x+30
525=7x
/7
x=75 m
length = 6x+30
6*75+30
480
The dimensions of pool 480m by 75m